the sum of the present ages of a father and son is 53 years. four years ago, the father's age was four times the age of the son. find their present ages.
Answer by ankor@dixie-net.com(22740)

the sum of the present ages of a father and son is 53 years.
f + s = 53
f = (53-s); use this from for substitution
:
four years ago, the father's age was four times the age of the son.
f - 4 = 4(s-4)
f - 4 = 4s - 16
f = 4s - 16 + 4
f = 4s - 12
replace f with (53-s)
53 - s = 4s - 12
53 + 12 = 4s + s
65 = 5s
s = 65/5
s = 13 is the son's age
I'll let you find the father's age, check your solutions in the 2nd statement